1. Find the kinetic energy of a photoelectron emitted from a rubidium (Rb) surface if the wavelength of the incident light is 400 nm (1 nm = 1·10^−9 m), and the binding (or threshold) energy of an electron is 3.6·10^−19J.

(A) 5.0·10−19 J

(B) 3.6·10−19J

(C) 1.4·10−19 J

(D) λthreshold > 400 nm, no emission of photoelectrons

(E) λthreshold < 400 nm, no emission of photoelectrons

Please show how you got the problem
Answer #1

Ans:

Let ‘E’ be the energy associated with the incident light

According to photoelectric effect,

E = BE + KE

Where, E = energy associated with the incident light = (h.c/ λ)

            h = plank’s constant = 6.626 × 10-34 m2 kg s-1

c = speed of light = 3 x 108 m/s

            λ = wavelength of incident light = 400 x 10-9 m

            BE = binding energy of electron = 3.6 x 10-19 J

            KE = kinetic energy of emitted electron

KE = E – BE = php95u8mz.png

KE = {(6.626 × 10-34 m2 kg s-1 x 3 x 108 m s-1)/ 400 x 10-9 m} - 3.6 x 10-19 J

{Note: 1 J = 1 kg m2s-2 }

KE = 1.3695 x 10-19 J

Kinetic energy of a photoelectron emitted from a rubidium (Rb) surface = 1.3695 x 10-19 J